# BARI-TH/25-772

### v1

@@ 3.3 $(g-2)_{e,\mu}$
The $Z^{\prime}$ contribution to the muon anomalous magnetic moment $a_{\mu}=\frac{(g-2)_{\mu}}{2}$ has been obtained in [1] using the general formulae in [53]:
@@ 4 Numerical analysis, results and correlations
As mentioned in Sec. 2, in the ABCD model various scenarios have been considered for the $Z^{\prime}$ couplings to fermions [1]. One of them, the scenario A, consists in assuming flavour violating couplings of $Z^{\prime}$ only with left-handed fermions, obtained setting to zero the parameters in (36)-(37). The present study is based on this minimal scenario.
[…]
@@
The allowed regions in the $(\epsilon_{1},\,\epsilon_{2})$ plane are displayed in Fig. 2 for the two values of $M_{Z^{\prime}}$. The regions are broader than those found in [1], a consequence of scanning $|V_{cb}|$ and $|V_{ub}|$ in the exclusive-inclusive ranges instead of choosing a preferred value. A remarkable result, shown in Fig. 3, is that the inclusive value $|V_{ub}|_{incl}$ in (69) is excluded by the $\Delta F=2$ constraints, which only allow for a smaller value. Provided with the regions of parameters, we can analyse the above listed processes.
@@ 4.1 $b\to s\ell_{1}^{+}\ell_{2}^{-}$
In SM where only $\ell_{1}=\ell_{2}$ is allowed, the values of the Wilson coefficients of the operators $O_{7}$ and $O_{9,10}$ in the low-energy Hamiltonian (38) read [55]:
@@
with $C_{7}^{\rm eff}$ defined in Sec. 3.1. Efforts have been spent in global analyses to identify the coefficients deviating from the SM value, and to which extent, to reproduce the experimental data [56, 57, 58, 59].
@@
In the scenario A of the ABCD model there are no contributions to the coefficients of the primed operators in (38). The $Z^{\prime}$ contribution to $C_{7}$ is negligible with respect to SM [54]. The contributions to $C_{9}$ and $C_{10}$ for $M_{Z^{\prime}}=1$ TeV and $3$ TeV are displayed in Figs. 4 and 5, where each point corresponds to a pair real/imaginary part of the coefficients computed for the same values of the parameters. The contributions to the real parts can be sizeable, up to ±$\pm 10\%$ of the SM values. Imaginary parts of ${\cal O}(10^{-4})$ are generated, originating from the phases of the CKM and PMNS matrices. The correlations between the real parts of the coefficients are displayed in the plots in Fig. 6 for $M_{Z^{\prime}}=1$ and $3$ TeV.
@@ 4.1.1 Exclusive lepton flavour conserving $b\to s\ell^{+}\ell^{-}$ modes
The simplest process induced by $b\to s\ell^{+}\ell^{-}$ is the purely leptonic $B_{s}$ decay, that we analyse using $f_{B_{s}}=230.3\pm 1.3$ MeV [5]. Fig. 7 shows the branching fractions ¯${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})$ and ¯${\bar{\cal B}}(B_{s}\to\tau^{+}\tau^{-})$ for $M_{Z^{\prime}}=1$ TeV and $3$ TeV. In SM we find ¯${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})=(3.65\pm 0.10)\times 10^{-9}\left(|\lambda_{ts}^{*}|/0.04\right)^{2}$ and ¯${\bar{\cal B}}(B_{s}\to\tau^{+}\tau^{-})=(7.70\pm 0.10)\times 10^{-7}\left(|\lambda_{ts}^{*}|/0.04\right)^{2}$. The experimental world average ¯${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})_{\rm exp}=(3.34\pm 0.27)\times 10^{-9}$ [24] is indicated in Fig. 7. The results in the ABCD model encompass the experimental range. Moreover, requiring that ¯${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})$ agrees with experiment, ¯${\bar{\cal B}}(B_{s}\to\tau^{+}\tau^{-})=(8.2\pm 0.5)\times 10^{-7}$ is predicted for $M_{Z^{\prime}}=1$ TeV, and ¯${\bar{\cal B}}(B_{s}\to\tau^{+}\tau^{-})=(8.2\pm 0.8)\times 10^{-7}$ for $M_{Z^{\prime}}=3$ TeV. The correlation between the two purely leptonic modes is shown in the figure, where each point corresponds to the branching fractions of $B_{s}\to\mu^{+}\mu^{-}$ and $B_{s}\to\tau^{+}\tau^{-}$ computed with the same set of parameters.
@@
For the lepton flavour conserving modes $B\to K^{*}\ell^{+}\ell^{-}$, with $\ell=\mu,\,\tau$, we use the local form factors determined in [60], see appendix A.
@@
Figure 7: Correlation between ¯${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})$ and ¯${\bar{\cal B}}(B_{s}\to\tau^{+}\tau^{-})$ for $M_{Z^{\prime}}=1$ TeV (left) and $M_{Z^{\prime}}=3$ TeV (right panel). The shaded region corresponds to the experimental world average for ¯${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})$[24].
@@
Figure 8: Correlation between ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\mu^{-})$ and ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\tau^{-})$ for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel). The shaded band corresponds to the experimental result ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\mu^{-})=(9.4\pm 0.5)\times 10^{-7}$ [24].
@@
Figure 8 shows the correlation between the branching fractions ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\mu^{-})$ and ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\tau^{-})$. The vertical band corresponds to the measurement ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\mu^{-})_{\rm exp}=(9.4\pm 0.5)\times 10^{-7}$ [24]. In the case of $B\to K^{*}\tau^{+}\tau^{-}$ only an upper limit is known: ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\tau^{-})<3.1\times 10^{-3}$. In the ABCD model the result for ${\cal B}(\bar{B}^{0}\to K^{*0}\mu^{+}\mu^{-})$ encompasses the experimental interval. Requiring that it lies in the measured range, the predictions ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\tau^{-})=(8.6\pm 1.3)\times 10^{-8}$ and ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\tau^{-})=(8.6\pm 0.9)\times 10^{-8}$ are obtained for $M_{Z^{\prime}}=1$ and $3$ TeV, respectively. For the SM we have: ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\mu^{-})=(9.6\pm 2.0)\times 10^{-7}\left(|\lambda_{ts}^{*}|/0.04\right)^{2}$ and ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\tau^{-})=(9.2\pm 2.0)\times 10^{-8}\left(|\lambda_{ts}^{*}|/0.04\right)^{2}$.
[…]
@@
Figure 11: Forward-backward asymmetry $A_{FB}(q^{2})$ (left) and $K^{*}$ polarization fraction $F_{L}(q^{2})$ (right plots) for $\bar{B}^{0}\to\bar{K}^{*0}(K\pi)\mu^{+}\mu^{-}$ (top panels) and $\bar{B}^{0}\to\bar{K}^{*0}(K\pi)\tau^{+}\tau^{-}$ (bottom panels). The dark continuous line is the SM result, the shaded region the ABCD result for $M_{Z^{\prime}}=1$ TeV. The black dots in the top panels correspond to the LHCb measurement [72].
@@
For the $P$-observables in (57) where tensions emerged [69, 70, 71, 72, 73, 74, 41], we display the most significant results in Fig. 12. The effects obtained in the ABCD model are more pronounced for $\tau$, namely for low $q^{2}$ the observable $P^{\prime}_{5}$ can differ from SM.
@@
Figure 12: Angular $P$-observables in Eq. (57) for $\bar{B}^{0}\to\bar{K}^{*0}(K\pi)\mu^{+}\mu^{-}$ (top panels) and $\bar{B}^{0}\to\bar{K}^{*0}(K\pi)\tau^{+}\tau^{-}$ (bottom panels). The black dots in the top panels correspond to LHCb measurement [72]. Same color code as in Fig. 11.
@@
For $B_{s}\to\phi\mu^{-}\mu^{+}$, using the form factors in [60] with parameters in Table 2, we find in SM: ¯${\bar{\cal{B}}}(B_{s}\to\phi\mu^{-}\mu^{+})=(11.8\pm 2.3)\times 10^{-7}\left(|\lambda_{ts}^{*}|/0.04\right)^{2}$, confirming that a higher value is foreseen for this rate with respect to the measured ¯${\bar{\cal{B}}}(B_{s}\to\phi\mu^{-}\mu^{+})_{\rm exp}=(8.4\pm 0.4)\times 10^{-7}$ [24]. Measurements are not available for the $\tau$ mode; we find ¯${\bar{\cal{B}}}(B_{s}\to\phi\tau^{-}\tau^{+})=(1.02\pm 0.20)\times 10^{-7}\left(|\lambda_{ts}^{*}|/0.04\right)^{2}$ in SM. Fig. 13 shows the correlation between the two LFC decays in the ABCD model: the plot for $M_{Z^{\prime}}=1$ TeV shows that the range for the $\tau$ mode is ¯${\bar{\cal{B}}}(B_{s}\to\phi\tau^{-}\tau^{+})=(0.9\pm 0.25)\times 10^{-7}$, for $M_{Z^{\prime}}=3$ TeV the range is ¯${\bar{\cal B}}(B_{s}\to\phi\tau^{-}\tau^{+})=(0.9\pm 0.2)\times 10^{-7}$. However, since ¯${\bar{\cal B}}(B_{s}\to\phi\mu^{-}\mu^{+})_{\rm exp}$ is not reproduced, a puzzle remains unsolved for this mode. If the measurement is confirmed, the first issue to scrutinise is the uncertainty of the $B_{s}\to\phi$ form factors.
@@
Figure 13: Correlation between ¯${\bar{\cal{B}}}(B_{s}\to\phi\mu^{+}\mu^{-})$ and ¯${\bar{\cal{B}}}(B_{s}\to\phi\tau^{+}\tau^{-})$ for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel). The shaded vertical band corresponds to the result ¯${\bar{\cal B}}(B_{s}\to\phi\mu^{+}\mu^{-})|_{exp}=(8.4\pm 0.4)\times 10^{-7}$ quoted in [24].
   LFV mode ×${\cal B}\times 10^{9}$ ${\cal B}^{(1)}\times 10^{9}$ ${\cal B}^{(2)}\times 10^{9}$ experiment
$B_{s}\to\tau^{+}\mu^{-}$ $0.00\div 2.10$ $0.00\div 2.10$ $0.00\div 1.60$ $<4.2\times 10^{-5}$ [24]
$\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-}$ $0.00\div 2.90$ $0.00\div 2.90$ $0.00\div 1.15$ $<1.0\times 10^{-5}$ [24]
$B_{s}\to\phi\tau^{+}\mu^{-}$ $0.00\div 3.43$ $0.00\div 2.60$ $<1.0\times 10^{-5}$[75]
$B_{s}\to\tau^{+}\mu^{-}$ $0.00\div 2.30$ $0.00\div 1.50$ $0.00\div 0.14$ $<4.2\times 10^{-5}$ [24]
$\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-}$ $0.00\div 3.10$ $0.00\div 2.00$ $0.00\div 0.20$ $<1.0\times 10^{-5}$ [24]
$B_{s}\to\phi\tau^{+}\mu^{-}$ $0.00\div 3.70$ $0.00\div 0.25$ $<1.0\times 10^{-5}$[75]

### v2

@@ 3.3 $(g-2)_{e,\mu}$
The $Z^{\prime}$ contribution to the muon anomalous magnetic moment $a_{\mu}=\frac{(g-2)_{\mu}}{2}$ has been obtained in [Aebischer:2019blw] using the general formulae in [Lindner:2016bgg]:
@@ 4 Numerical analysis, results and correlations
As mentioned in Sec. 2, in the ABCD model various scenarios have been considered for the $Z^{\prime}$ couplings to fermions [Aebischer:2019blw]. One of them, the scenario A, consists in assuming flavour violating couplings of $Z^{\prime}$ only with left-handed fermions, obtained setting to zero the parameters in (36)-(37). The present study is based on this minimal scenario.
The list of the parameters consists of the four entries in (35) and of the parameters of the CKM and PMNS matrices. The coupling $g_{Z^{\prime}}$ is varied in the range $g_{Z^{\prime}}\in[0.01,\,0.1]$, and two cases are considered for the $Z^{\prime}$ mass, $M_{Z^{\prime}}=1$ TeV and $M_{Z^{\prime}}=3$ TeV. The four independent parameters of the CKM matrix are taken as $V_{us},\,\gamma,\,|V_{cb}|$ and $|V_{ub}|$, see appendix C. The strategy adopted in the present study consists in varying $|V_{cb}|$ and $|V_{ub}|$ in the ranges determined by the exclusive and inclusive measurements,
@@
The allowed regions in the $(\epsilon_{1},\,\epsilon_{2})$ plane are displayed in Fig. 2 for the two values of $M_{Z^{\prime}}$. The regions are broader than those found in [Aebischer:2019blw], a consequence of scanning $|V_{cb}|$ and $|V_{ub}|$ in the exclusive-inclusive ranges instead of choosing a preferred value.
A remarkable result, shown in Fig. 3, is that the scan over the parameter space subject to the $\Delta F=2$ constraints does not yield any configuration with $|V_{ub}|$ equal to the inclusive determination. The largest obtained value is $|V_{ub}|=3.92\times 10^{-3}$ (within the resolution set by the scan step), which indicates a preference for smaller values. This agrees with the conclusions of the discussion in [Blanke:2018cya] on the sensitivity of $\Delta F=2$ observables to the precise values of the CKM parameters.
After having imposed the $\Delta F=2$ constraints, we first look for deviations in lepton-flavour-conserving (LFC) $B$ and $B_{s}$ decays and for the correlations with the corresponding lepton-flavour-violating (LFV) modes. The LFC processes provide non-trivial constraints on the LFV ones.
We then study the correlations with purely leptonic LFV processes, in particular with $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$, $\mu^{-}\to e^{-}\gamma$, $\mu^{-}\to e^{-}e^{+}e^{-}$ and $\mu^{-}\to e^{-}$ conversion in nuclei, obtaining that the experimental upper bounds on such purely leptonic channels play a hierarchical role in constraining the branching ratios of LFV $B$ and $B_{s}$ decays. The bound on $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$ does not impose significant restrictions, whereas the other three channels set progressively more stringent limits. The correlations with $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$, $\mu^{-}\to e^{-}\gamma$ allow branching ratios for the LFV beauty meson decays of ${\cal O}(10^{-9})$. Including constraints from $\mu^{-}\to e^{-}e^{+}e^{-}$ and $\mu^{-}\to e^{-}$ conversion in nuclei, the branching fractions are pushed down to ${\cal O}(10^{-11})$ and ${\cal O}(10^{-13})$, as discussed below.
@@ 4.1 $b\to s\ell_{1}^{+}\ell_{2}^{-}$
In SM where only $\ell_{1}=\ell_{2}$ is allowed, the values of the Wilson coefficients of the operators $O_{7}$ and $O_{9,10}$ in the low-energy Hamiltonian (38) read [EOSAuthors:2021xpv]:
@@
with $C_{7}^{\rm eff}$ defined in Sec. 3.1. Efforts have been spent in global analyses to identify the coefficients deviating from the SM value, and to which extent, to reproduce the experimental data [Altmannshofer:2014rta, Capdevila:2017bsm, Alguero:2021anc, Alguero:2023jeh].
@@
In the scenario A of the ABCD model there are no contributions to the coefficients of the primed operators in (38). The $Z^{\prime}$ contribution to $C_{7}$ is negligible with respect to SM [Buras:2012jb]. The contributions to $C_{9}$ and $C_{10}$ for $M_{Z^{\prime}}=1$ TeV and $3$ TeV are displayed in Figs. 4 and 5, where each point corresponds to a pair real/imaginary part of the coefficients computed for the same values of the parameters. The contributions to the real parts can be sizeable, up to $\pm 10\%$ of the SM values. Imaginary parts of ${\cal O}(10^{-4})$ are generated, originating from the phases of the CKM and PMNS matrices. The correlations between the real parts of the coefficients are displayed in the plots in Fig. 6 for $M_{Z^{\prime}}=1$ and $3$ TeV.
@@ 4.1.1 Exclusive lepton flavour conserving $b\to s\ell^{+}\ell^{-}$ modes
The simplest process induced by $b\to s\ell^{+}\ell^{-}$ is the purely leptonic $B_{s}$ decay, that we analyse using $f_{B_{s}}=230.3\pm 1.3$ MeV [HeavyFlavorAveragingGroupHFLAV:2024ctg]. Fig. 7 shows the branching fractions ${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})$ and ${\bar{\cal B}}(B_{s}\to\tau^{+}\tau^{-})$ for $M_{Z^{\prime}}=1$ TeV and $3$ TeV. In SM we find ${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})=(3.65\pm 0.10)\times 10^{-9}\left(|\lambda_{ts}^{*}|/0.04\right)^{2}$ and ${\bar{\cal B}}(B_{s}\to\tau^{+}\tau^{-})=(7.70\pm 0.10)\times 10^{-7}\left(|\lambda_{ts}^{*}|/0.04\right)^{2}$. The experimental world average ${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})_{\rm exp}=(3.34\pm 0.27)\times 10^{-9}$ [Navas:PDG] is indicated in Fig. 7. The results in the ABCD model encompass the experimental range. Moreover, requiring that ${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})$ agrees with experiment, ${\bar{\cal B}}(B_{s}\to\tau^{+}\tau^{-})=(8.2\pm 0.5)\times 10^{-7}$ is predicted for $M_{Z^{\prime}}=1$ TeV, and ${\bar{\cal B}}(B_{s}\to\tau^{+}\tau^{-})=(8.2\pm 0.8)\times 10^{-7}$ for $M_{Z^{\prime}}=3$ TeV. The correlation between the two purely leptonic modes is shown in the figure, where each point corresponds to the branching fractions of $B_{s}\to\mu^{+}\mu^{-}$ and $B_{s}\to\tau^{+}\tau^{-}$ computed with the same set of parameters. Fig. 7 shows that when ${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})$ is predicted within its experimental range, the branching ratio ${\bar{\cal B}}(B_{s}\to\tau^{+}\tau^{-})$ is also constrained. This feature emerges from the numerical scan, and can be understood from the analytic structure of the amplitudes in scenario A. Indeed, Eq. (49) shows that both branching ratios are proportional to $|C_{10}|^{2}$,
${\bar{\cal B}}(B_{s}\to\ell^{+}\ell^{-})=K_{\ell}\times|C_{10}|^{2}=K_{\ell}\times\left|C_{10}^{SM}+(C_{10}^{NP})_{\ell\ell}\right|^{2}\,\,\,$, (76)
with $K_{\ell}$ a numerical factor different for $\ell=\mu$ and $\ell=\tau$. From Eq. (75) one observes that $(C_{10}^{NP})_{\mu\mu}$ and $(C_{10}^{NP})_{\tau\tau}$ differ for the term $\Delta_{A}^{\ell\ell}$ that depends on the scanned parameters. Using Eq. (28) one finds
$(C_{10}^{NP})_{\tau\tau}\simeq(C_{10}^{NP})_{\mu\mu}+f(g_{Z},\,\epsilon_{1},\,\epsilon_{2})\,\,$, (77)
where $f$ depends on the scanned parameters. Hence, one writes
${\bar{\cal B}}(B_{s}\to\tau^{+}\tau^{-})\simeq\frac{\lambda^{1/2}(m_{B_{s}}^{2},m_{\tau}^{2},m_{\tau}^{2})}{\lambda^{1/2}(m_{B_{s}}^{2},m_{\mu}^{2},m_{\mu}^{2})}\left[{\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})+{\tilde{f}}(g_{Z},\,\epsilon_{1},\,\epsilon_{2})\right]\,\,$, (78)
where ${\tilde{f}}(g_{Z},\,\epsilon_{1},\,\epsilon_{2})$ denotes the terms taking into account the function $f$ in Eq. (77). This relation shows that the experimental bound on ${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})$ also limits ${\bar{\cal B}}(B_{s}\to\tau^{+}\tau^{-})$.
@@
For the lepton flavour conserving modes $B\to K^{*}\ell^{+}\ell^{-}$, with $\ell=\mu,\,\tau$, we use the local form factors determined in [Bharucha:2015bzk] and described in appendix A.
@@
Figure 7: Correlation between ${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})$ and ${\bar{\cal B}}(B_{s}\to\tau^{+}\tau^{-})$ for $M_{Z^{\prime}}=1$ TeV (left) and $M_{Z^{\prime}}=3$ TeV (right panel). The shaded region corresponds to the experimental world average for ${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})$[Navas:PDG].
@@
Figure 8: Correlation between ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\mu^{-})$ and ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\tau^{-})$ for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel). The shaded band corresponds to the experimental result ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\mu^{-})=(9.4\pm 0.5)\times 10^{-7}$ [Navas:PDG].
@@
Figure 8 shows the correlation between the branching fractions ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\mu^{-})$ and ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\tau^{-})$. The vertical band corresponds to the measurement ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\mu^{-})_{\rm exp}=(9.4\pm 0.5)\times 10^{-7}$ [Navas:PDG]. In the case of $B\to K^{*}\tau^{+}\tau^{-}$ only an upper limit is known: ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\tau^{-})<3.1\times 10^{-3}$. In the ABCD model the result for ${\cal B}(\bar{B}^{0}\to K^{*0}\mu^{+}\mu^{-})$ encompasses the experimental interval. Requiring that it lies in the measured range, the predictions ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\tau^{-})=(8.6\pm 1.3)\times 10^{-8}$ and ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\tau^{-})=(8.6\pm 0.9)\times 10^{-8}$ are obtained for $M_{Z^{\prime}}=1$ and $3$ TeV, respectively. For the SM we have: ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\mu^{-})=(9.6\pm 2.0)\times 10^{-7}\left(|\lambda_{ts}^{*}|/0.04\right)^{2}$ and ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\tau^{-})=(9.2\pm 2.0)\times 10^{-8}\left(|\lambda_{ts}^{*}|/0.04\right)^{2}$.
The $B\to K^{*}(K\pi)\mu^{+}\mu^{-}$ angular distribution has been experimentally studied in details [BaBar:2006tnv, BaBar:2015wkg, Belle:2009zue, Belle:2016fev, CDF:2011tds, CMS:2015bcy, CMS:2017rzx, ATLAS:2018gqc], finding a deviation from SM predictions in some angular $P$-observables defined in (57) [LHCb:2013zuf, LHCb:2013ghj, LHCb:2015svh, LHCb:2020lmf, LHCb:2020gog, LHCb:2023gpo, LHCb:2024onj]. In the analysis of the ABCD model predictions, to make the comparison with SM more transparent, we define the rescaled angular coefficient functions ${\tilde{I}}_{i}=I_{i}/|\lambda_{ts}^{*}|^{2}$, with $i=1s,\dots 6s$, and display them in Figs. 9, 10 for $\ell=\mu$ and $\ell=\tau$, setting $M_{Z^{\prime}}=1$ TeV (for $M_{Z^{\prime}}=3$ TeV the results do not change sensibly). The functions $I_{6c,\,7,\,8,\,9}$ vanish in SM; in the ABCD model they are about $10^{-6}$ times smaller than all others, therefore we do not display them. In the figures, the dark continuous line is the SM result for the central values of the form factor parameters, the light shaded band the ABCD result. Modest deviations with respect to SM are possible, they are more pronounced in the case of $\tau$.
@@
Figure 11: Forward-backward asymmetry $A_{FB}(q^{2})$ (left) and $K^{*}$ polarization fraction $F_{L}(q^{2})$ (right plots) for $\bar{B}^{0}\to\bar{K}^{*0}(K\pi)\mu^{+}\mu^{-}$ (top panels) and $\bar{B}^{0}\to\bar{K}^{*0}(K\pi)\tau^{+}\tau^{-}$ (bottom panels). The dark continuous line is the SM result, the shaded region the ABCD result for $M_{Z^{\prime}}=1$ TeV. The black dots in the top panels correspond to the LHCb measurement [LHCb:2020lmf].
@@
For the $P$-observables in (57) where tensions emerged [LHCb:2013zuf, LHCb:2013ghj, LHCb:2015svh, LHCb:2020lmf, LHCb:2020gog, LHCb:2023gpo, LHCb:2024onj], we display the most significant results in Fig. 12. The effects obtained in the ABCD model are more pronounced for $\tau$, namely for low $q^{2}$ the observable $P^{\prime}_{5}$ can differ from SM.
@@
Figure 12: Angular $P$-observables in Eq. (57) for $\bar{B}^{0}\to\bar{K}^{*0}(K\pi)\mu^{+}\mu^{-}$ (top panels) and $\bar{B}^{0}\to\bar{K}^{*0}(K\pi)\tau^{+}\tau^{-}$ (bottom panels). The black dots in the top panels correspond to LHCb measurement [LHCb:2020lmf]. Same color code as in Fig. 11.
@@
For $B_{s}\to\phi\mu^{-}\mu^{+}$, using the form factors in [Bharucha:2015bzk] with parameters in Table 3, we find in SM: ${\bar{\cal{B}}}(B_{s}\to\phi\mu^{-}\mu^{+})=(11.8\pm 2.3)\times 10^{-7}\left(|\lambda_{ts}^{*}|/0.04\right)^{2}$, confirming that a higher value is foreseen for this rate with respect to the measured ${\bar{\cal{B}}}(B_{s}\to\phi\mu^{-}\mu^{+})_{\rm exp}=(8.4\pm 0.4)\times 10^{-7}$ [Navas:PDG]. Measurements are not available for the $\tau$ mode: we find ${\bar{\cal{B}}}(B_{s}\to\phi\tau^{-}\tau^{+})=(1.02\pm 0.20)\times 10^{-7}\left(|\lambda_{ts}^{*}|/0.04\right)^{2}$ in SM. Fig. 13 shows the correlation between the two LFC decays in the ABCD model: the plot for $M_{Z^{\prime}}=1$ TeV shows that the range for the $\tau$ mode is ${\bar{\cal{B}}}(B_{s}\to\phi\tau^{-}\tau^{+})=(0.9\pm 0.25)\times 10^{-7}$, for $M_{Z^{\prime}}=3$ TeV the range is ${\bar{\cal B}}(B_{s}\to\phi\tau^{-}\tau^{+})=(0.9\pm 0.2)\times 10^{-7}$. However, since ${\bar{\cal B}}(B_{s}\to\phi\mu^{-}\mu^{+})_{\rm exp}$ is not reproduced, a puzzle remains unsolved for this mode. If the measurement is confirmed, the first issue to scrutinise is the uncertainty of the $B_{s}\to\phi$ form factors.
@@
Figure 13: Correlation between ${\bar{\cal{B}}}(B_{s}\to\phi\mu^{+}\mu^{-})$ and ${\bar{\cal{B}}}(B_{s}\to\phi\tau^{+}\tau^{-})$ for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel). The shaded vertical band corresponds to the result ${\bar{\cal B}}(B_{s}\to\phi\mu^{+}\mu^{-})|_{exp}=(8.4\pm 0.4)\times 10^{-7}$ quoted in [Navas:PDG].
   LFV mode ${\cal B}^{(1)}\times 10^{9}$ ${\cal B}^{(2)}\times 10^{9}$ ${\cal B}^{(3)}\times 10^{11}$ ${\cal B}^{(4)}\times 10^{13}$ experiment
$B_{s}\to\tau^{+}\mu^{-}$ $00\div 2.10 0.0$$0\div 1.60 0.0$$0\div 1.20 0.0$$00\div 9.20$ $4.2\times 10^{-5}$ [Navas:PDG]
$\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-}$ $00\div 2.90 0.0$$0\div 1.15 0.0$$0\div 1.60 0.0$$00\div 5.10$ $1.0\times 10^{-5}$ [Navas:PDG]
$B_{s}\to\phi\tau^{+}\mu^{-}$ $00\div 3.43 0.0$$0\div 2.60 0.0$$0\div 1.95 0.0$$00\div 6.10$ $1.0\times 10^{-5}$[LHCb:2024wve]
$B_{s}\to\tau^{+}\mu^{-}$ $0\div 2.30 0.00÷0$$\div 0.14 0.0$$0\div 1.10 0.0$$00\div 1.05$ $4.2\times 10^{-5}$ [Navas:PDG]
$\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-}$ $0\div 3.10 0.00÷0$$\div 0.20 0.0$$0\div 1.53 0.0$$00\div 1.42$ $1.0\times 10^{-5}$ [Navas:PDG]
$B_{s}\to\phi\tau^{+}\mu^{-}$ $0\div 3.70 0.00÷0$$\div 0.25 0.0$$0\div 1.80 0.0$$00\div 1.70$ $1.0\times 10^{-5}$[LHCb:2024wve]